# Understanding Deep Representation Learning via Layerwise Feature Compression and Discrimination

### v3

@@ B.1 Properties of Weight Matrices
$\sigma_{1}^{2}$ $=\sum_{i=1}^{p+q}\sigma_{i}^{2}\left\|\bm{V}_{l+1}^{(i)^{\top}}\bm{U}_{l}^{(1)}\bm{x}\right\|^{2}\leq\sigma_{1}^{2}\left\|\bm{V}_{l+1}^{(1)^{\top}}\bm{U}_{l}^{(1)}\bm{x}\right\|^{2}+\sigma_{2}^{2}\sum_{i=2}^{p+q}\left\|\bm{V}_{l+1}^{(i)^{\top}}\bm{U}_{l}^{(1)}\bm{x}\right\|^{2}$
$=\sigma_{1}^{2}\left\|\bm{V}_{l+1}^{(1)^{\top}}\bm{U}_{l}^{(1)}\bm{x}\right\|^{2}+\sigma_{2}^{2}\left(1-\left\|\bm{V}_{l+1}^{(1)^{\top}}\bm{U}_{l}^{(1)}\bm{x}\right\|^{2}\right)=\sigma_{2}^{2}+\left(\sigma_{1}^{2}-\sigma_{2}^{2}\right)\left\|\bm{V}_{l+1}^{(1)^{\top}}\bm{U}_{l}^{(1)}\bm{x}\right\|^{2},$
where the inequality follows from $\sigma_{1}>\sigma_{2}>\dots>\sigma_{p+q}$, and the second equality uses $\bm{V}_{l+1}\in\mathcal{O}^{d}$. This, together with $\sigma_{1}>\sigma_{2}>0$, implies $\left\|\bm{V}_{l+1}^{(1)^{\top}}\bm{U}_{l}^{(1)}\bm{x}\right\|^{2}=1$. Using this and $\|\bm{x}\|=1$, we have $\bm{U}_{l}^{(1)}=\bm{V}_{l+1}^{(1)}\bm{Q}_{l,1}$ for some $\bm{Q}_{l,1}\in\mathcal{O}^{h_{1}}$. By repeatedly applying the same argument to $\bm{U}_{l}^{(i)}$ for all $i\geq 2$, we prove (56).
@@
$\|\bm{V}_{L,1}\bm{\Sigma}^{2}_{L,1}\bm{V}_{L,1}^{\top}-\bm{U}_{L-1,1}\bm{\Sigma}^{2}_{L-1,1}\bm{U}_{L-1,1}^{\top}-\bm{U}_{L-1,2}\bm{\Sigma}_{L-1,2}^{2}\bm{U}_{L-1,2}^{\top}\|_{F}\leq\delta.$
Using this, $\|\bm{U}^{\top}\bm{A}\bm{U}\|_{F}\leq\|\bm{A}\|_{F}$ for any $\bm{U}\in\mathcal{O}^{d\times(d-K)}$, we further obtain
$\|\bm{U}_{L-1,2}^{\top}\bm{V}_{L,1}\bm{\Sigma}^{2}_{L,1}\bm{V}_{L,1}^{\top}\bm{U}_{L-1,2}-\bm{\Sigma}_{L-1,2}^{2}\|_{F}\leq\delta.$
@@
$\|\bm{U}_{L-1,2}^{\top}\bm{V}_{L,1}\bm{\Sigma}^{2}_{L,1}\bm{V}_{L,1}^{\top}\bm{U}_{L-1,2}\|_{F}\leq 2\delta.$
Using Lemma 1 with the fact that $\bm{\Sigma}_{L,1}\bm{V}_{L,1}^{\top}\bm{U}_{L-1,2}$ is of rank $K$, we further have
$\|\bm{\Sigma}_{L,1}\bm{V}_{L,1}^{\top}\bm{U}_{L-1,2}\|_{F}^{2}\leq 2\delta\sqrt{K}.$ (59)
Noting that $\|\bm{\Sigma}_{L,1}\bm{V}_{L,1}^{\top}\bm{U}_{L-1,2}\|_{F}\geq\sigma_{\min}(\bm{\Sigma}_{L,1})\|\bm{V}_{L,1}^{\top}\bm{U}_{L-1,2}\|_{F}$, we have
$\|\bm{V}_{L,1}^{\top}\bm{U}_{L-1,2}\|_{F}\leq\frac{\|\bm{\Sigma}_{L,1}\bm{V}_{L,1}^{\top}\bm{U}_{L-1,2}\|_{F}}{\sigma_{\min}(\bm{\Sigma}_{L,1})}\leq\frac{\sqrt{2\delta}\sqrt[4]{K}}{\left(n/2\right)^{1/2L}}\leq\frac{2\sqrt{\delta}\sqrt[4]{K}}{n^{1/2L}},$ (60)
@@
$\sigma_{\min}^{2}(\bm{V}_{L,1}^{\top}\bm{U}_{L-1,1})$ $=\min_{\|\bm{x}\|=1}\|\bm{U}_{L-1,1}^{\top}\bm{V}_{L,1}\bm{x}\|^{2}=\min_{\|\bm{x}\|=1}\bm{x}^{\top}\bm{V}_{L,1}^{\top}\bm{U}_{L-1,1}\bm{U}_{L-1,1}^{\top}\bm{V}_{L,1}\bm{x}$
$=\min_{\|\bm{x}\|=1}\bm{x}^{\top}\bm{V}_{L,1}^{\top}\left(\bm{I}-\bm{U}_{L-1,2}\bm{U}_{L-1,2}^{\top}\right)\bm{V}_{L,1}\bm{x}=1-\max_{\|\bm{x}\|=1}\|\bm{U}_{L-1,2}^{\top}\bm{V}_{L,1}\bm{x}\|^{2}$
$=1-\sigma_{\max}^{2}(\bm{U}_{L-1,2}^{\top}\bm{V}_{L,1})\geq 1-\|\bm{U}_{L-1,2}^{\top}\bm{V}_{L,1}\|_{F}^{2},$
@@
$\sigma_{\min}(\bm{V}_{L,1}^{\top}\bm{U}_{L-1,1})\geq\sqrt{1-\|\bm{U}_{L-1,2}^{\top}\bm{V}_{L,1}\|_{F}^{2}}\geq 1-\|\bm{U}_{L-1,2}^{\top}\bm{V}_{L,1}\|_{F}\geq 1-\frac{2\sqrt{\delta}\sqrt[4]{K}}{n^{1/2L}}.$
B.2 Analysis of Progressive Within-Class Compression
@@
$\bm{W}_{l:1}=\bm{U}_{l,1}\bm{O}_{l,1}\tilde{\bm{\Sigma}}^{l}_{1}\bm{V}_{1,1}^{\top}+\bm{U}_{l,2}\bm{O}_{l,2}\tilde{\bm{\Sigma}}^{l}_{2}\bm{V}_{1,2}^{\top}=\bm{U}_{l,1}\tilde{\bm{\Sigma}}^{l}_{1}\bm{O}_{l,1}\bm{V}_{1,1}^{\top}+\bm{U}_{l,2}\tilde{\bm{\Sigma}}^{l}_{2}\bm{O}_{l,2}\bm{V}_{1,2}^{\top},$ (65)
@@
To bound $\mathrm{Tr}(\bm{\Sigma}_{W}^{l})$ and $\mathrm{Tr}(\bm{\Sigma}_{B}^{l})$ across layers, we present a lemma that establishes a relationship between the magnitudes of $\bm{V}_{1,1}^{\top}\bm{\Delta}_{W}$ and $\bm{V}_{1,2}^{\top}\bm{\Delta}_{W}$, as well as a relationship between the magnitudes of $\bm{V}_{1,1}^{\top}\bm{\Delta}_{B}$ and $\bm{V}_{1,2}^{\top}\bm{\Delta}_{B}$ based on the above setup.
Lemma 9.
Suppose that Assumptions 1 and 2 hold with
@@
Proof.
(i) It follows from (10) that $\bm{W}_{L:1}\bm{X}=\bm{Y}$. This, together with $\bm{Y}=\bm{I}_{K}\otimes\bm{1}_{n}$ and (27), implies $\bm{W}_{L:1}\bm{x}_{k,i}=\bm{e}_{k}$ and $\bm{W}_{L:1}\bar{\bm{x}}_{k}=\bm{e}_{k}$ for all $i\in[n]$ and $k\in[K]$. It follows from this and (61) that
@@
where the first inequality uses Lemma 2, and the last inequality follows from (45) in Lemma 7 and (57) in Lemma 8. This, together with (B.2), yields
[…]
@@
[…]
@@
be a compact eigenvalue decomposition of $\bm{\Delta}_{1}$, where $\bm{U}_{\Delta_{1}}\in\mathcal{O}^{d\times K}$ and $\bm{\Sigma}_{\Delta_{1}}\in\mathbb{R}^{K\times K}$. Using this and Lemma 3 with $\bm{\Delta}_{1}=\bm{V}_{1,1}\bm{A}_{1}\bm{A}_{1}^{T}\bm{V}_{1,1}^{T}$ by (B.2), we obtain
@@
be a compact eigenvalue decomposition of $\bm{W}_{L:1}^{T}\bm{W}_{L:1}$, where $\bm{U}_{1}\in\mathbb{R}^{d\times K}$ and $\bm{\Sigma}_{1}\in\mathbb{R}^{K\times K}$. According to (9) and (29), we have $\sigma_{K}(\bm{W}_{L:1})\geq\sqrt{n/(1-\theta)}$. This, together with (B.2), (78), and Davis-Kahan Theorem (see, e.g., [104, Theorem V.3.6]), yields
@@
[…]
@@
where the second inequality follows from (81), (B.2), and (84), the last inequality uses $\theta\leq 1/4$ and $n\geq 16$. This, together with $\|\bm{\Delta}_{B}\|_{F}^{2}=\|\bm{V}_{1,1}^{T}\bm{\Delta}_{B}\|_{F}^{2}+\|\bm{V}_{1,2}^{T}\bm{\Delta}_{B}\|_{F}^{2}$, yields
@@
which, together with $\theta\leq 1/4$, implies (68). ∎
@@
Theorem 2.
Suppose that Assumptions 1 and 2 hold with $\delta,\varepsilon,\rho$ satisfying (13). Then, we have
@@
Proof.
For ease of exposition, let $\bm{v}_{i}\in\mathbb{R}^{d}$ denote the $i$-th column of $\bm{V}_{1}\in\mathcal{O}^{d}$. Therefore, we write
@@
where the equality follows from (88), the second inequality uses (12), and the last inequality is due to (B.2). This, together with $\theta\leq 1/4$, (30), (61), (87), and (B.2), yields
@@
Using (B.2) and (B.2) again, we have for all $l\in[L-2]$,
@@
B.3 Analysis of Progressive Between-Class Discrimination
@@
where $\bm{z}_{k,i}^{l}$ and $\bar{\bm{x}}_{k}$ are defined in (4) and (27), respectively. Moreover, recall from (25) that $\bm{V}_{1,1}$ is the right singular matrix of $\bm{W}_{1}$. Using the results in Sections B.1 and B.2, we are ready to derive the explicit form of $\bm{V}_{1,1}\bar{\bm{x}}_{k}$ and bound the norm of $\bm{\mu}_{k}^{l}$.
Lemma 10.
Suppose that Assumptions 1 and 2 hold with $\delta$ and $\varepsilon$ satisfying (13). Then, the following statements hold:
@@
Proof.
According to (9), we have $\bm{W}_{L:1}\bm{X}=\bm{Y}$. This, together with $\bm{Y}=\bm{I}_{K}\otimes\bm{1}_{n}^{T}$ and (27), implies
@@
[…]
@@
where the first inequality uses the triangular inequality and the fact that $\bm{O}$ and $\tilde{\bm{\Sigma}}_{1}$ can be interchanged, the second inequality follows from (45) and (B.3), and the last inequality follows from (13). This, together with (106), implies (100). ∎
@@
Theorem 3.
Suppose that Assumptions 1 and 2 hold with $\delta,\varepsilon,\rho$ satisfying (13). Then, we have
@@
Proof.
According to (65) and (98), we compute for all $k\neq k^{\prime}$,
@@
where the second inequality follows from (B.3) and (113). Next, we compute
@@
where the last inequality follows from (13). These, together with (B.3), yield
@@
where the second inequality follows from $\|\bar{\bm{x}}_{k}\|^{2}\leq\|\sum_{i=1}^{n}\bm{x}_{k,i}/n\|^{2}\leq 1/n+\theta/N$ due to Assumption 1. This, together with (100) in Lemma 10, (110), and (B.3), yields for all $k,l$,
@@
This, together with (8), yields (109). ∎
B.4 Proof of 1
Based on 2 and 3, we can directly obtain 1.
B.5 Proof of 1

### v4

@@ A.1 Properties of Weight Matrices
$\sigma_{1}^{2}$ $=\sum_{i=1}^{p+q}\sigma_{i}^{2}\left\|\bm{V}_{l+1}^{(i)^{T}}\bm{U}_{l}^{(1)}\bm{x}\right\|^{2}\leq\sigma_{1}^{2}\left\|\bm{V}_{l+1}^{(1)^{T}}\bm{U}_{l}^{(1)}\bm{x}\right\|^{2}+\sigma_{2}^{2}\sum_{i=2}^{p+q}\left\|\bm{V}_{l+1}^{(i)^{T}}\bm{U}_{l}^{(1)}\bm{x}\right\|^{2}$
$=\sigma_{1}^{2}\left\|\bm{V}_{l+1}^{(1)^{T}}\bm{U}_{l}^{(1)}\bm{x}\right\|^{2}+\sigma_{2}^{2}\left(1-\left\|\bm{V}_{l+1}^{(1)^{T}}\bm{U}_{l}^{(1)}\bm{x}\right\|^{2}\right)=\sigma_{2}^{2}+\left(\sigma_{1}^{2}-\sigma_{2}^{2}\right)\left\|\bm{V}_{l+1}^{(1)^{T}}\bm{U}_{l}^{(1)}\bm{x}\right\|^{2},$
where the inequality follows from $\sigma_{1}>\sigma_{2}>\dots>\sigma_{p+q}$, and the second equality uses $\bm{V}_{l+1}\in\mathcal{O}^{d}$. This, together with $\sigma_{1}>\sigma_{2}>0$, implies $\left\|\bm{V}_{l+1}^{(1)^{T}}\bm{U}_{l}^{(1)}\bm{x}\right\|^{2}=1$. Using this and $\|\bm{x}\|=1$, we have $\bm{U}_{l}^{(1)}=\bm{V}_{l+1}^{(1)}\bm{Q}_{l,1}$ for some $\bm{Q}_{l,1}\in\mathcal{O}^{h_{1}}$. By repeatedly applying the same argument to $\bm{U}_{l}^{(i)}$ for all $i\geq 2$, we prove (49).
@@
$\|\bm{V}_{L,1}\bm{\Sigma}^{2}_{L,1}\bm{V}_{L,1}^{T}-\bm{U}_{L-1,1}\bm{\Sigma}^{2}_{L-1,1}\bm{U}_{L-1,1}^{T}-\bm{U}_{L-1,2}\bm{\Sigma}_{L-1,2}^{2}\bm{U}_{L-1,2}^{T}\|_{F}\leq\delta.$
Using this, $\|\bm{U}^{T}\bm{A}\bm{U}\|_{F}\leq\|\bm{A}\|_{F}$ for any $\bm{U}\in\mathcal{O}^{d\times(d-K)}$, we further obtain
$\|\bm{U}_{L-1,2}^{T}\bm{V}_{L,1}\bm{\Sigma}^{2}_{L,1}\bm{V}_{L,1}^{T}\bm{U}_{L-1,2}-\bm{\Sigma}_{L-1,2}^{2}\|_{F}\leq\delta.$
@@
$\|\bm{U}_{L-1,2}^{T}\bm{V}_{L,1}\bm{\Sigma}^{2}_{L,1}\bm{V}_{L,1}^{T}\bm{U}_{L-1,2}\|_{F}\leq 2\delta.$
Using Lemma 14 with the fact that $\bm{\Sigma}_{L,1}\bm{V}_{L,1}^{T}\bm{U}_{L-1,2}$ is of rank $K$, we further have
$\|\bm{\Sigma}_{L,1}\bm{V}_{L,1}^{T}\bm{U}_{L-1,2}\|_{F}^{2}\leq 2\delta\sqrt{K}.$ (52)
Noting that $\|\bm{\Sigma}_{L,1}\bm{V}_{L,1}^{T}\bm{U}_{L-1,2}\|_{F}\geq\sigma_{\min}(\bm{\Sigma}_{L,1})\|\bm{V}_{L,1}^{T}\bm{U}_{L-1,2}\|_{F}$, we have
$\|\bm{V}_{L,1}^{T}\bm{U}_{L-1,2}\|_{F}\leq\frac{\|\bm{\Sigma}_{L,1}\bm{V}_{L,1}^{T}\bm{U}_{L-1,2}\|_{F}}{\sigma_{\min}(\bm{\Sigma}_{L,1})}\leq\frac{\sqrt{2\delta}\sqrt[4]{K}}{\left(n/2\right)^{1/2L}}\leq\frac{2\sqrt{\delta}\sqrt[4]{K}}{n^{1/2L}},$ (53)
@@
$\sigma_{\min}^{2}(\bm{V}_{L,1}^{T}\bm{U}_{L-1,1})$ $=\min_{\|\bm{x}\|=1}\|\bm{U}_{L-1,1}^{T}\bm{V}_{L,1}\bm{x}\|^{2}=\min_{\|\bm{x}\|=1}\bm{x}^{T}\bm{V}_{L,1}^{T}\bm{U}_{L-1,1}\bm{U}_{L-1,1}^{T}\bm{V}_{L,1}\bm{x}$
$=\min_{\|\bm{x}\|=1}\bm{x}^{T}\bm{V}_{L,1}^{T}\left(\bm{I}-\bm{U}_{L-1,2}\bm{U}_{L-1,2}^{T}\right)\bm{V}_{L,1}\bm{x}=1-\max_{\|\bm{x}\|=1}\|\bm{U}_{L-1,2}^{T}\bm{V}_{L,1}\bm{x}\|^{2}$
$=1-\sigma_{\max}^{2}(\bm{U}_{L-1,2}^{T}\bm{V}_{L,1})\geq 1-\|\bm{U}_{L-1,2}^{T}\bm{V}_{L,1}\|_{F}^{2},$
@@
$\sigma_{\min}(\bm{V}_{L,1}^{T}\bm{U}_{L-1,1})\geq\sqrt{1-\|\bm{U}_{L-1,2}^{T}\bm{V}_{L,1}\|_{F}^{2}}\geq 1-\|\bm{U}_{L-1,2}^{T}\bm{V}_{L,1}\|_{F}\geq 1-\frac{2\sqrt{\delta}\sqrt[4]{K}}{n^{1/2L}}.$
A.2 Analysis of Progressive Within-Class Compression
@@
$\bm{W}_{l:1}=\bm{U}_{l,1}\bm{O}_{l,1}\tilde{\bm{\Sigma}}^{l}_{1}\bm{V}_{1,1}^{T}+\bm{U}_{l,2}\bm{O}_{l,2}\tilde{\bm{\Sigma}}^{l}_{2}\bm{V}_{1,2}^{T}=\bm{U}_{l,1}\tilde{\bm{\Sigma}}^{l}_{1}\bm{O}_{l,1}\bm{V}_{1,1}^{T}+\bm{U}_{l,2}\tilde{\bm{\Sigma}}^{l}_{2}\bm{O}_{l,2}\bm{V}_{1,2}^{T},$ (58)
@@
To bound $\mathrm{Tr}(\bm{\Sigma}_{W}^{l})$ and $\mathrm{Tr}(\bm{\Sigma}_{B}^{l})$ across layers, we present a lemma that establishes a relationship between the magnitudes of $\bm{V}_{1,1}^{T}\bm{\Delta}_{W}$ and $\bm{V}_{1,2}^{T}\bm{\Delta}_{W}$, as well as a relationship between the magnitudes of $\bm{V}_{1,1}^{T}\bm{\Delta}_{B}$ and $\bm{V}_{1,2}^{T}\bm{\Delta}_{B}$ based on the above setup.
Lemma 10
Suppose that Assumptions 2 and 3 hold with
@@
Proof (i) It follows from (10) that $\bm{W}_{L:1}\bm{X}=\bm{Y}$. This, together with $\bm{Y}=\bm{I}_{K}\otimes\bm{1}_{n}$ and (23), implies $\bm{W}_{L:1}\bm{x}_{k,i}=\bm{e}_{k}$ and $\bm{W}_{L:1}\bar{\bm{x}}_{k}=\bm{e}_{k}$ for all $i\in[n]$ and $k\in[K]$. It follows from this and (54) that
@@
where the first inequality uses Lemma 15, and the last inequality follows from (39) in Lemma 8 and (50) in Lemma 9. This, together with (A.2), yields
$\left\|\bm{V}_{1,1}^{T}\bm{\Delta}_{W}\right\|_{F}\leq\frac{\frac{\sqrt{2\delta}\sqrt[4]{K}}{\left(n/2\right)^{1/2L}}\varepsilon^{L-1}}{\left(1-\frac{\sqrt{2\delta}\sqrt[4]{K}}{\left(n/2\right)^{1/2L}}\right)\left(\frac{n}{2}\right)^{\frac{L-1}{2L}}}\left\|\bm{V}_{1,2}^{T}\bm{\Delta}_{W}\right\|_{F}\leq\frac{2\sqrt{2\delta}\sqrt[4]{K}\varepsilon^{L-1}}{\sqrt{n}}\left\|\bm{V}_{1,2}^{T}\bm{\Delta}_{W}\right\|_{F},$
@@
$\leq\left(\sqrt{2n}\right)^{\frac{L+1}{L}}\varepsilon^{L-1}\frac{\sqrt{2\delta}\sqrt[4]{K}}{(n/2)^{1/2L}}\leq 2^{1+\frac{1}{L}}\varepsilon^{L-1}\sqrt{n\delta}\sqrt[4]{K},$ (67)
@@
be a compact eigenvalue decomposition of $\bm{\Delta}_{1}$, where $\bm{U}_{\Delta_{1}}\in\mathcal{O}^{d\times K}$ and $\bm{\Sigma}_{\Delta_{1}}\in\mathbb{R}^{K\times K}$. Using this and Lemma 16 with $\bm{\Delta}_{1}=\bm{V}_{1,1}\bm{A}_{1}\bm{A}_{1}^{T}\bm{V}_{1,1}^{T}$ by (A.2), we obtain
@@
be a compact eigenvalue decomposition of $\bm{W}_{L:1}^{T}\bm{W}_{L:1}$, where $\bm{U}_{1}\in\mathbb{R}^{d\times K}$ and $\bm{\Sigma}_{1}\in\mathbb{R}^{K\times K}$. According to (9) and (25), we have $\sigma_{K}(\bm{W}_{L:1})\geq\sqrt{n/(1-\theta)}$. This, together with (A.2), (71), and Davis-Kahan Theorem (see, e.g., (stewart1990matrix, Theorem V.3.6)), yields
@@
where the second inequality uses (A.2) and (68), and the last inequality follows from $4\varepsilon^{L-1}\sqrt{\delta}\sqrt[4]{K}\leq 4\sqrt{\delta}\sqrt[4]{K}\leq(6-4\sqrt{2})\sqrt{n}$ due to $\varepsilon\leq 1$ and (59). Using this and (72) yields
@@
where the second inequality follows from (74), (A.2), and (77), the last inequality uses $\theta\leq 1/4$ and $n\geq 16$. This, together with $\|\bm{\Delta}_{B}\|_{F}^{2}=\|\bm{V}_{1,1}^{T}\bm{\Delta}_{B}\|_{F}^{2}+\|\bm{V}_{1,2}^{T}\bm{\Delta}_{B}\|_{F}^{2}$, yields
@@
which, together with $\theta\leq 1/4$, implies (61).
@@
Theorem 11
Suppose that Assumptions 2 and 3 hold with $\delta,\varepsilon,\rho$ satisfying (13). Then, we have
@@
Proof For ease of exposition, let $\bm{v}_{i}\in\mathbb{R}^{d}$ denote the $i$-th column of $\bm{V}_{1}\in\mathcal{O}^{d}$. Therefore, we write
@@
where the equality follows from (81), the second inequality uses (12), and the last inequality is due to (A.2). This, together with $\theta\leq 1/4$, (26), (54), (80), and (A.2), yields
@@
Using (A.2) and (A.2) again, we have for all $l\in[L-2]$,
@@
A.3 Analysis of Progressive Between-Class Discrimination
@@
where $\bm{z}_{k,i}^{l}$ and $\bar{\bm{x}}_{k}$ are defined in (4) and (23), respectively. Moreover, recall from (21) that $\bm{V}_{1,1}$ is the right singular matrix of $\bm{W}_{1}$. Using the results in Sections A.1 and A.2, we are ready to derive the explicit form of $\bm{V}_{1,1}\bar{\bm{x}}_{k}$ and bound the norm of $\bm{\mu}_{k}^{l}$.
Lemma 12
Suppose that Assumptions 2 and 3 hold with $\delta$ and $\varepsilon$ satisfying (13). Then, the following statements hold:
@@
Proof According to (9), we have $\bm{W}_{L:1}\bm{X}=\bm{Y}$. This, together with $\bm{Y}=\bm{I}_{K}\otimes\bm{1}_{n}^{T}$ and (23), implies
@@
$\leq\frac{\sqrt{2\delta}\sqrt[4]{K}}{n^{1/2L}}\left(\sqrt{2n}\right)^{\frac{L-1}{L}}\|\bar{\bm{x}}_{k}\|+\frac{\sqrt{2\delta}\sqrt[4]{K}\varepsilon^{L-1}}{n^{1/2L}}\|\bar{\bm{x}}_{k}\|\leq\frac{5\sqrt{\delta}\sqrt[4]{K}}{4n^{1/2L}}\left(\sqrt{2n}\right)^{\frac{L-1}{L}}\left(\frac{1}{\sqrt{n}}+\sqrt{\frac{\theta}{nK}}\right),$
@@
where the first inequality uses the triangular inequality and the fact that $\bm{O}$ and $\tilde{\bm{\Sigma}}_{1}$ can be interchanged, the second inequality follows from (39) and (A.3), and the last inequality follows from (13). This, together with (99), implies (93).
@@
Theorem 13
Suppose that Assumptions 2 and 3 hold with $\delta,\varepsilon,\rho$ satisfying (13). Then, we have
@@
Proof According to (58) and (91), we compute for all $k\neq k^{\prime}$,
@@
where the second inequality follows from (A.3) and (106). Next, we compute
@@
where the last inequality follows from (13). These, together with (A.3), yield
@@
where the second inequality follows from $\|\bar{\bm{x}}_{k}\|^{2}\leq\|\sum_{i=1}^{n}\bm{x}_{k,i}/n\|^{2}\leq 1/n+\theta/N$ due to Assumption 2. This, together with (93) in Lemma 12, (103), and (A.3), yields for all $k,l$,
@@
This, together with (8), yields (102).
A.4 Proof of Theorem 4
Based on Theorem 11 and Theorem 13, we can directly obtain Theorem 4.
A.5 Proof of Proposition 5